# Interpretable Pre-Trained Transformers for Heart Time-Series Data

### v1

@@ Abstract
Decoder-only transformers are the backbone of the popular generative pre-trained transformer (GPT) series of large language models. In this work, we apply the same framework to periodic heart time-series data to create two pre-trained general purpose cardiac models, namely PPG-PT and ECG-PT. We demonstrate that both such pre-trained models are fully interpretable. This is achieved firstly through aggregate attention maps which show that the model focuses on similar points in previous cardiac cycles in order to make predictions and gradually broadens its attention in deeper layers. Next, tokens with the same value, that occur at different distinct points in the ECG and PPG cycle, form separate clusters in high dimensional space based on their phase as they propagate through the transformer blocks. Finally, we highlight that individual attention heads respond to specific physiologically relevent features, such as the dicrotic notch in PPG and the P-wave in ECG. It is also demonstrated that these pre-trained models can be easily fine-tuned for tasks such as classification of atrial fibrillation. In this specific example, the fine-tuning took 11 minutes of computer time, and achieved a leave-one-subject-out AUCs of 0.99 and 0.93 for ECG and PPG respectively. Importantly, these fine-tuned models are also fully explainable, with attention shifting to regions in the context that are strongly indicative of atrial fibrillation.
@@ 1 Introduction
The generative pretrained transformer (GPT) [1] [2] is a large language model (LLM) which forms the basis of the chat-GPT models [3]. It consists of layers of decoder-only transformers [4] that are trained to predict the next token (words, sub-words, or characters) using all previous tokens provided to the model as context. The attention mechanism is masked such that tokens inputted to the model can only communicate with past tokens and thus a context window of $N$ length effectively provides $N$ different training examples, whereby the model tries to predict the same context window shifted by one token into the future. The whole premise of the GPT model is that in order to accurately predict the next token, across a large and diverse array of texts, a model must gain an efficient and general comprehension of these texts. Given that language is a way that we summarise and communicate the world around us, this efficient and general understanding allows the pre-trained model to be fine tuned for a myriad of complex tasks [3].
@@
The premise of this paper is that we can use this same GPT idea, i.e. stacks of decoder-only transformers trained to predict the next token, and apply it to physiological time series instead of language. With the hypothesis that if our models gain an efficient and general understanding of these physiological time series, such as photoplethysmography (PPG) or electrocardiography (ECG), then these large models will be simple to fine-tune for downstream tasks such as the classification of heart conditions. The ECG signal refers to the monitoring of the electrical activity of the heart through external non-invasive electrodes that measure the potential difference across the heart [5][6]. In this work, we use only single-lead ECG. The photoplethysmography (PPG) signal refers to the non-invasive measurement of blood volume [7]. The PPG works by shining light through tissue with an LED, which is absorbed by the blood, and then either measuring the transmitted light or reflected light with a corresponding photo-diode. The PPG is thus able to measure changes in blood volume related to pulse and even breathing [8]. In PPG, for example, in order to accurately predict the next time point the model must gain knowledge on underlying several phenomena, such as the subtle changes in heart rate and pulse amplitude due to breathing and blood pressure. We can therefore hypothesise that a large model that is good at predicting the next token of PPG will be straightforward to fine-tune for specific PPG related tasks.
@@
As GPT and other large language models rapidly gain capability and thus decision making power in our lives, it is of crucial importance that LLMs do not remain black boxes. The black box nature of deep learning models is also an immediate problem in healthcare, where decisions made by artificial intelligence have the potential to directly impact our health and even our lives. It is therefore critical that models are developed which give clear insights and explanations into why decisions are made. A logical way to interpret the operation of transformer-based LLMs is to look at the attention weights of the individual transformer blocks. This attention mechanism provides information on which tokens a model looks at in order to make a decision, and thus it is common to use heat maps that visualise this attention as an avenue for interpretability [9]. Language is complex, and there are often several pathways to achieve the same result. In many instances this makes it difficult to decipher attention, and means that in the context of LLMs it cannot always be relied upon for explainability [10]. However, both the PPG and ECG signals are far less complicated than language, and downstream tasks, such as screening for heart conditions, offer more constraints on the possible pathways to achieve a correct diagnosis. It is therefore a core focus of this paper, not just to provide generalised models for PPG and ECG related tasks, but to provide models that are fully interpretable in their operation so that they may be safely used in a healthcare setting.
@@
To this end, we firstly present two pretrained decoder-only time series transformer models for use with heart signals, namely PPG-PT and ECG-PT. Furthermore, we demonstrate that, through i) aggregate attention of different transformer layers, ii) changes in cosine similarity between core PPG and ECG features in embedding space upon propagation through transformer layers, and iii) the analysis of the attention weights of individual attention heads in the final transformer block, that both the PPG-PT and ECG-PT models behave exactly as expected for the task of predicting the next time series point in PPG and ECG. Finally, we demonstrate how these models can be easily fined tuned to detect atrial fibrillation (AF), a common type of abnormal heart rhythm. The changes in attention from the task of next token prediction, to the task of classification of AF, can be used to further explain why a model has made a decision. Clinicians can thus be provided with both a classification and the reasoning behind it, allowing for a collaborative relationship between a clinician and AI in diagnostics and treatment.
@@ 2.1 Tokenisation and Embedding
Both the ECG and PPG signals are periodic, and in most applications we can ignore any mean offset. With this in mind, it is simple to divide each PPG and ECG signal into a range of finite tokens, forming a vocabulary from which each signal can be constructed. To tokenise the PPG signals, they were resampled to 50Hz and each 10 second context window (500 samples) was scaled to between 100 and 0. This signal was then rounded to give integer values, yielding a total vocabulary size of 101 tokens. The 10 second context window was chosen so as to be long enough to preserve low frequency variations in the PPG (such as respiratory variations), but not too long as to run into memory issues when training. For the ECG, the process was so, apart from resampling to 100Hz and thus using a context window that corresponded to 5 seconds instead of 10 seconds. A higher sampling frequency was required in the ECG tokenisation in order to preserve the high frequency components of the QRS complex. In GPT models, common sequences of token values can be combined during tokenisation, in order to allow for a longer context length for the same number of tokens. This method of tokenisation was not implemented in the PPG-PT and ECG-PT models as further extending the length of the context window was not necessary.
Upon tokenisation, each token was embedded with a token embedding table and a position embedding table. The dimensions of the token embedding were the vocabulary size times the embedding dimension ($d_{model}$), giving a $d_{model}$ dimensional vector for every possible token in the vocabulary. The dimensions of the position embedding table were the maximum context length times $d_{model}$, giving a $d_{model}$-dimensional vector for every possible position in the context window. In our models, both of these embeddings are learnt as the model trains. In the original transformer paper, the positional embeddings utilised sine and cosines functions of different frequencies [4]. The token and positional embeddings are added together, which means that the attention mechanism in the subsequent transformer blocks has information on both the specific token and the position of the token in the context.
@@ 2.2 Multi-Head Masked Self Attention and the Transformer Block
Once tokens have been embedded with a token embedding and positional embedding that are added together, thus allowing the model to use information on both the position of the token in the context window (length $N$) and the value of the token, they are fed to transformer blocks in their high dimensional vector form. In multi-head attention, the embedding space (of size $d_{model}$) is divided up into lower dimensional spaces of equal size ($d_{k}$), before the results of attention are concatenated after the attention mechanism to reconstruct the original embedding space size ($d_{model}$) [4].
The attention mechanism works by allowing tokens to communicate with each other and thus update each others information. Each attention head transforms tokens into keys ($K$), using a linear layer which compresses the tokens from $d_{model}$ dimensions to $d_{k}$ dimensions. It also separately transforms tokens into queries ($Q$), which again use another linear layer to compress the tokens from $d_{model}$ dimensions to $d_{k}$ dimensions. One can think of queries ($Q$) as each token broadcasting what it is looking for, and keys ($K$) as each token broadcasting what it contains. The dot product is then taken between both the queries and the keys, such that if there is a match between the information a token is looking for in $Q$ and the information a token emits in $K$, there will be a spike in the corresponding attention weights $QK^{T}$. This dot product is also scaled by the $\sqrt{d_{k}}$, and the soft max of each row is taken (resulting in attention rows that sum to 1). This scaled-dot product results in an $N\times N$ matrix of attention weights, which show how tokens communicate with each other, for all tokens in the context window. In masked self attention, the mechanism used to train GPT models, the upper triangle of this weight matrix is set to zero. This means that tokens can only communicate with themselves or with past tokens in the sequence, and allows for efficient training by allowing for N separate training examples with an input block of N token length. During training of the model, dropout (the random zeroing of weights) is also applied to the weight matrix. The final step to masked self attention is to multiply the weight matrix with a value matrix. The value matrix, $V$, is formed from another linear layer that compresses the tokens from $d_{model}$ dimensions to $d_{k}$ dimensions. A combination of all of these steps produces the following attention function [4].
@@ 2.3 Architecture
Our models were developed in PyTorch [11] and adapted from a tutorial given by Andrej Karpathy [12] titled "Let’s build GPT: from scratch, in code, spelled out". For both the photoplethysmography pre-trained transformer (PPG-PT) and electrocardiography pre-trained transformer (ECG-PT) we used an embedding dimension ($d_{model}$) of 64. The original GPT paper used an embedding dimension of 768 [1], but given that PPG and ECG are assumed to be far less complex than language, 64 was found to be sufficient in our case. We used 8 transformer blocks, each with 8 attention heads, as described in Section 2.2 of this paper and the original transformer paper [4]. It was found during early training that bigger models performed better, so we trained the largest model possible, with the trade-offs of context length ($N=$ 500 samples) and batch size (64), in order to utilise almost all of the 12GB of RAM on our NVIDIA RTX A2000 GPU. This architecture resulted in 443,493 trainable parameters. In the areas where dropout was applied (the feedforward subsection of the transformer and the attention weight matrix), dropout was set to 0.2.
2.4 Training
2.4.1 Pre-Training
@@
Three datasets were used for training the PPG-PT base model: 1) Capnobase “respiratory benchmark" dataset [13], which consists of high quality ECG and PPG recorded over 42 subjects of 8 minutes each; 2) BIDMC [^1] The BIDMC data set is made available under the ODC Attribution license, and is available at https://physionet.org/content/bidmc/1.0.0/ dataset [14] which consists of PPG and ECG from 53 subjects for 8 minutes each; 3) the “cuffless blood pressure dataset" [15], a subset of MIMIC II [16], consisting of 12,000 recordings of PPG signals of varied quality in a hospital setting. The combination of all of these datasets resulted in over 128 million total tokens for training.
@@
For the training of the ECG-PT base model, we used subsets of the "PhysioNet Computing in Cardiology Challenge 2020" dataset [17] [16], which consists of tens of thousands of 12-lead ECG recordings, across tens of thousands of patients in a hospital setting. From each of these recordings, we extracted 10 second examples of Lead I ECG. Within this dataset are a diverse range of cardiac abnormalities as well as many healthy subjects, as the dataset was originally constructed for the identification of different heart conditions. Once tokenised, this dataset resulted in over 42 million tokens for training.
@@
For each model, training data was split into 90% training and 10% validation datasets. The data was not shuffled to ensure that validation data primarily consisted of unseen subjects. The PPG-PT model was trained over 500,000 iterations with a batch size of 64, and the ECG-PT model was trained over 560,000 iterations with the same batch size. After every 2000 iterations, the models were evaluated for a further 200 iterations to measure validation loss. In both cases the learning rate was set as $3\times 10^{-4}$ within the AdamW optimiser. To train PPG-PT, it took just over 5 days on an RTX A2000 12GB GPU, and to train ECG-PT it took almost 6 days. Optimisation loss was measured using cross entropy, by mapping the $N\times 101$ model outputs (logits) where the $101$ represents all possible tokens, to the target tokens which were the next token values for all points in the input. The final validation loss of the PPG-PT model was 1.2, corresponding to roughly a third of prediction being perfect. The final validation loss of the ECG-PT model was 1.8 (1 in 6 predictions are perfect). The higher validation loss for ECG-PT was likely due to the high proportion of abnormal heart rhythms in the training and validation datasets.
2.4.2 Atrial Fibrillation Fine-Tuning
@@
The final linear layer, which converts the output from $d_{model}$ dimensions to the dimensions of the vocabulary size was replaced with a linear layer which instead converts the output from $d_{model}$ dimensions to 1 dimension. This new output then was passed through a sigmoid activation function to scale the values between 0 and 1, and the final value was used to classify the signal as either healthy (0) or atrial fibrillation (1). This same conversion of the final linear layer of the model could be used for any suitable classification or regression task, by simply scaling the number of output dimensions to be in line with the specific task. During training, just the new linear layer and the final transformer block were trained, whilst all other layers were frozen. In this example, the learning rate was maintained at $3\times 10^{-4}$ within the AdamW optimiser, and a binary cross entropy loss function was used.
@@
To train and evaluate the fine tuning of the model for classification of atrial fibrillation (AF), the "MIMIC PERform AF" dataset [18] was used, which is a subset of MIMIC III dataset [19]. This dataset contains 20 minutes of continuous PPG and ECG recordings from each of 35 subjects, 19 of whom have AF and 16 of whom did not have AF. The signals, originally recorded with a sample frequency of 125Hz, were band-pass filtered to between 1 and 15Hz and then downsampled to 50Hz in the case of the PPG, and band-pass filtered between 1 and 45Hz and downsampled to 100Hz in the case of the ECG. A window length of 500 samples was maintained, corresponding to 10 seconds of PPG (50Hz) and 5 seconds of ECG (100Hz), and sliding windows were extracted from the data and tokenised with a shift of 50 samples each time, in order to artificially inflate the volume of data for the model to train on. In both cases of the ECG and PPG, the models were trained in a leave-one-subject-out fashion by training on 34 subjects and testing on 1, over 1000 iterations each with a batch size of 128. Each model took 11 minutes to fine tune on an RTX A2000 GPU, which is a fraction of the over 5 days that it took to train each of the base models.

### v2

@@ Abstract
Decoder-only transformers are the backbone of the popular generative pre-trained transformer (GPT) series of large language models. In this work, we employ this framework to the analysis of clinical heart time-series data, to create two pre-trained general purpose cardiac models, termed PPG-PT and ECG-PT. We place a special emphasis on making both such pre-trained models fully interpretable. This is achieved firstly through aggregate attention maps which show that, in order to make predictions, the model focuses on similar points in previous cardiac cycles and gradually broadens its attention in deeper layers. Next, we show that tokens with the same value, which occur at different distinct points in the electrocardiography (ECG) and photoplethysmography (PPG) cycle, form separate clusters in high dimensional space. The clusters form according to phase, as the tokens propagate through the transformer blocks. Finally, we highlight that individual attention heads respond to specific physiologically relevent features, such as the dicrotic notch in PPG and the P-wave in ECG. It is also demonstrated that these pre-trained models are straightforward to fine-tune for tasks such as classification of atrial fibrillation (AF), and beat detection in photoplethysmography. For the example of AF, the fine-tuning took 11 minutes of computer time, and achieved the respective leave-one-subject-out AUCs of 0.99 and 0.93 for ECG and PPG within the MIMIC Perform AF dataset. In addition, the fine-tuned beat detector achieved a state-of-the-art F1 score of 98%, as well as uniquely providing a beat confidence level which acts as a signal quality estimator. Importantly, the fine-tuned models for AF screening are also fully explainable, with attention shifting to regions in the context that are strongly indicative of atrial fibrillation.
@@ 1 Introduction
The generative pretrained transformer (GPT) [1] [2] is a large language model (LLM) which forms the basis of the chat-GPT models [3]. It consists of layers of decoder-only transformers [4] that are trained to predict the next token (words, sub-words, or characters), using all previous tokens provided to the model as context. The attention mechanism is masked, such that tokens inputted to the model can only communicate with past tokens and thus a context window of length $N$ effectively provides $N$ different training examples, whereby the model tries to predict the same context window shifted by one token into the future. The whole premise of the GPT model is that in order to accurately predict the next token, across a large and diverse array of texts, a model must gain an efficient and general comprehension of these texts. Given that language is a way that we summarise and communicate the world around us, this efficient and general understanding allows the pre-trained model to be fine tuned for a myriad of complex tasks [3].
@@
The assertion of this paper is that we can leverage the GPT mechanism, i.e. stacks of decoder-only transformers trained to predict the next token, and develop it for physiological time series. Our hypothesis is that if our models gain an efficient and general understanding of these physiological time series, such as photoplethysmography (PPG) or electrocardiography (ECG), then these large models will be simple to fine-tune for downstream tasks such as the classification of heart conditions. The ECG signal refers to the monitoring of the electrical activity of the heart through external non-invasive electrodes that measure the potential difference across the heart [5][6]. In this work, we use only single-lead ECG. The photoplethysmography (PPG) signal refers to the non-invasive measurement of blood volume [7]. The PPG operates by shining light through tissue with an LED, which is absorbed by the blood, and then either measuring the transmitted light or reflected light with a corresponding photo-diode. The PPG is thus able to measure changes in blood volume related to pulse and even breathing [8]. In PPG, for example, in order to accurately predict the next time point the model must gain knowledge on several underlying phenomena, such as the subtle changes in heart rate and pulse amplitude due to breathing and blood pressure. This leads us to hypothesise that a large model that is good at predicting the next token of PPG will be straightforward to fine-tune for specific PPG related tasks.
@@
As GPT and other large language models are rapidly gaining capability, and thus decision making power, it is of crucial importance that LLMs do not remain black boxes. Indeed, the black box nature of deep learning models is an immediate problem in healthcare, where decisions made by artificial intelligence have the potential to directly impact our health and even our lives. It is therefore critical that models are developed which give clear insights and explanations into why decisions are made. A natural way to interpret the operation of transformer-based LLMs is to take a close examination of the attention weights of the individual transformer blocks. This attention mechanism provides information on which tokens a model looks at in order to make a decision, and thus it is common to use heat maps that visualise this attention as an avenue for interpretability [9]. Language is complex, and there are often several pathways to achieve the same result. In many instances this makes it difficult to decipher attention, which means that in the context of LLMs attention cannot always be relied upon for explainability [10]. However, both the PPG and ECG signals are far less complicated than language, and downstream tasks, such as screening for heart conditions, offer more constraints on the possible pathways to achieve a correct diagnosis. It is therefore a core focus of this paper not just to provide generalised transformer models for PPG and ECG related tasks, but to provide models that are fully interpretable so that they may be safely used in a healthcare setting.
@@
To this end, we firstly develop two pre-trained decoder-only time series transformer models for use with heart signals, namely PPG-PT and ECG-PT. The pre-trained models and example code are provided at https://github.com/harryjdavies/HeartGPT. Furthermore, we demonstrate that, through i) aggregate attention of different transformer layers, ii) changes in cosine similarity between core PPG and ECG features in the embedding space upon propagation through transformer layers, and iii) the analysis of the attention weights of individual attention heads in the final transformer block, that both the PPG-PT and ECG-PT models behave exactly as expected for the task of predicting the next time series point in PPG and ECG. Finally, we demonstrate how these models can be fine-tuned to detect atrial fibrillation (AF), a common type of abnormal heart rhythm. The changes in attention from the task of next token prediction to the task of classification of AF can be used to further explain the reason why a model has made a decision. This makes it possible to provide clinicians with both accurate classification and the reasoning behind it, thus allowing for a conjoint operation between a clinician and AI in diagnostics and treatment.
@@ 2.1 Tokenisation and Embedding
Both the ECG and PPG signals are periodic, and in most applications we can ignore any mean offset. With this in mind, we can divide each PPG and ECG signal into a range of finite tokens, forming a vocabulary from which each signal can be constructed. To tokenise the PPG signals, they were resampled to 50Hz and each 10 second context window (500 samples) was scaled to between 100 and 0. The resulting signal was then rounded to give integer values, yielding a total vocabulary size of 101 tokens. The 10 second context window was chosen so as to be long enough to preserve low frequency variations in the PPG (such as respiratory variations), but not too long as to run into memory issues when training. For the ECG, the process was similar, apart from resampling to 100Hz and thus using a context window that corresponded to 5 seconds instead of 10 seconds. A higher sampling frequency was required in the ECG tokenisation in order to preserve the high frequency components of the QRS complex. In GPT models, common sequences of token values can be combined during tokenisation, in order to allow for a longer context length for the same number of tokens. This method of tokenisation was not implemented in the PPG-PT and ECG-PT models, as further extending the length of the context window was not necessary.
Upon tokenisation, each token was embedded with a token embedding table and a position embedding table. The dimensions of the token embedding were the vocabulary size times the embedding dimension ($d_{model}$), giving a $d_{model}$ dimensional vector for every possible token in the vocabulary. The dimensions of the position embedding table were the maximum context length times $d_{model}$, giving a $d_{model}$-dimensional vector for every possible position in the context window. In our models, both of these embeddings are learnt as the model trains. In the original transformer paper, the positional embeddings utilised sine and cosine functions of different frequencies [4]. The token and positional embeddings were added together, which means that the attention mechanism in the subsequent transformer blocks had information on both the specific token and the position of the token in the context.
@@ 2.2 Multi-Head Masked Self Attention and the Transformer Block
Once tokens have been embedded with a token embedding and positional embedding which are added together, thus allowing the model to use information on both the position of the token in the context window (length $N$) and the value of the token, these are fed to transformer blocks in their high dimensional vector form. In multi-head attention, the embedding space (of size $d_{model}$) is divided into lower dimensional spaces of equal size ($d_{k}$), before the results of attention are concatenated after the attention mechanism to reconstruct the original embedding space size ($d_{model}$) [4].
The attention mechanism operates by allowing tokens to communicate with one another and thus update each other’s information. Each attention head transforms tokens into keys ($K$), using a linear layer which compresses the tokens from $d_{model}$ dimensions to $d_{k}$ dimensions. It also separately transforms tokens into queries ($Q$), which again use another linear layer to compress the tokens from $d_{model}$ dimensions to $d_{k}$ dimensions. One can think of queries ($Q$) as each token broadcasting what it is looking for, and keys ($K$) as each token broadcasting what it contains. The dot product is then taken between both the queries and the keys, such that if there is a match between the information a token is looking for in $Q$ and the information a token emits in $K$, there will be a spike in the corresponding attention weights $QK^{T}$. This dot product is next scaled by $\sqrt{d_{k}}$, and the SoftMax of each row is taken (resulting in attention rows that sum to 1). This scaled dot product results in an $N\times N$ matrix of attention weights, which convey how tokens communicate with each other, for all tokens in the context window. In masked self attention, the mechanism used to train GPT models, the upper triangle of this weight matrix is set to zero. This means that tokens can only communicate with themselves or with past tokens in the sequence, and allows for efficient training by allowing for N separate training examples with an input block of N token length. During training of the model, dropout (the random zeroing of weights) is also applied to the weight matrix. The final step to the masked self attention process is to multiply the weight matrix with a value matrix. The value matrix, $V$, is formed from another linear layer that compresses the tokens from $d_{model}$ dimensions to $d_{k}$ dimensions. A combination of all of these steps produces the following attention function [4]
@@ 2.3 Architecture
Our models were developed in PyTorch [11] and adapted from a tutorial by Andrej Karpathy [12] titled “Let’s build GPT: from scratch, in code, spelled out”. For both the photoplethysmography pre-trained transformer (PPG-PT) and electrocardiography pre-trained transformer (ECG-PT) we used an embedding dimension ($d_{model}$) of 64. The original GPT paper used an embedding dimension of 768 [1], but given that PPG and ECG are far less complex than language, 64 was found to be sufficient in our case. We used 8 transformer blocks, each with 8 attention heads, as described in Section 2.2 of this paper and the original transformer paper [4]. It was found during early training that bigger models performed better, so we trained the largest model possible, with the trade-offs of context length ($N=$ 500 samples) and batch size (64), in order to utilise almost all of the 12GB of RAM on our NVIDIA RTX A2000 GPU. This architecture resulted in 443,493 trainable parameters. In the areas where dropout was applied (the feedforward subsection of the transformer and the attention weight matrix), dropout was set to 0.2.
2.4 Pre-Training
@@
Three datasets were used for training the PPG-PT base model: 1) Capnobase “respiratory benchmark” dataset [13], which consists of high quality ECG and PPG recorded over 42 subjects for 8 minutes each; 2) BIDMC [^1] The BIDMC data set is made available under the ODC Attribution license, and is available at https://physionet.org/content/bidmc/1.0.0/ dataset [14] which consists of PPG and ECG from 53 subjects for 8 minutes each; 3) the “cuffless blood pressure dataset” [15], a subset of MIMIC II [16], consisting of 12,000 recordings of PPG signals of varied quality, in a hospital setting. The combination of all of these datasets resulted in over 128 million total tokens for training.
@@
For the training of the ECG-PT base model, we used subsets of the “PhysioNet Computing in Cardiology Challenge 2020” dataset [17] [16], which consists of tens of thousands of 12-lead ECG recordings, across tens of thousands of patients in a hospital setting. From each of these recordings, we extracted 10-second examples of Lead I ECG. This dataset comprises a diverse range of cardiac abnormalities as well as many healthy subjects, as the dataset was originally constructed for the identification of different heart conditions. Once tokenised, this dataset resulted in over 42 million tokens for training.
@@
For each model, training data was split into 90% training and 10% validation datasets. The data was not shuffled to ensure that validation data primarily consisted of unseen subjects. The PPG-PT model was trained over 500,000 iterations with a batch size of 64, and the ECG-PT model was trained over 560,000 iterations with the same batch size. After every 2,000 iterations, the models were evaluated over a further 200 iterations to measure validation loss. In both cases, the learning rate was set as $3\times 10^{-4}$ within the AdamW optimiser. The training of PPG-PT took just over 5 days on an RTX A2000 12GB GPU, while the training of ECG-PT took almost 6 days. Optimisation loss was measured using cross entropy, by mapping the $N\times 101$ model outputs (logits) where the $101$ represents all possible tokens, to the target tokens which were the next token values for all points in the input. The final validation loss of the PPG-PT model was 1.2, corresponding to roughly a third of prediction being perfect. The final validation loss of the ECG-PT model was 1.8 (1 in 6 predictions are perfect). The higher validation loss for ECG-PT was likely due to the high proportion of abnormal heart rhythms in the training and validation datasets.
2.5 Fine-Tuning
2.5.1 Atrial Fibrillation
@@
During the fine-tuning, the final linear layer, which converts the output from $d_{model}$ dimensions to the dimensions of the vocabulary size, was replaced with a linear layer which instead converts the output from $d_{model}$ dimensions to 1 dimension. This new output then was passed through a sigmoid activation function to scale the values between 0 and 1; the final value was then used to classify the signal as either healthy (0) or atrial fibrillation (1). A binary cross entropy loss criterion was then used to map the output value to the true class. This same conversion of the final linear layer of the model could be used for any suitable classification or regression task, by simply scaling the number of output dimensions to be in line with the specific task. During training, just the new linear layer and the final transformer block were trained, whilst all other layers were frozen. In this example, the learning rate was maintained at $3\times 10^{-4}$ within the AdamW optimiser.
@@
To train and evaluate the fine-tuning of the model for classification of atrial fibrillation (AF), the “MIMIC PERform AF” dataset [18] was used, which is a subset of MIMIC III dataset [19]. This dataset contains 20 minutes of continuous PPG and ECG recordings from each of 35 subjects, of whom 19 had AF and 16 did not have AF. These signals, originally recorded at a sample frequency of 125Hz, were band-pass filtered to between 1 and 15Hz and then downsampled to 50Hz in the case of the PPG, and band-pass filtered between 1 and 45Hz and downsampled to 100Hz in the case of the ECG. A window length of 500 samples was maintained, corresponding to 10 seconds of PPG (50Hz) and 5 seconds of ECG (100Hz), and sliding windows were extracted from the data and tokenised with a shift of 50 samples each time, in order to artificially inflate the volume of data for the model training. In both the cases of ECG and PPG, the models were fine-tuned in a leave-one-subject-out fashion by training on 34 subjects and testing on 1, over 1,000 iterations and with a batch size of 128. Each model took 11 minutes to fine tune on an RTX A2000 GPU, which is a fraction of the over 5 days that it took to train each of the base models.
2.5.2 PPG Beat Detection
As was the case with fine-tuning to detect atrial fibrillation, the final linear layer was again replaced with a linear layer which converted the output from $d_{model}$ dimensions to 1 dimension. However, instead of using just the final value in the context as was the case with AF, all values from the context were used. This 1-dimensional array was passed through a Sigmoid activation function, to map to values of 0 (no peak) or 1 (peak) for all tokens in the context. Only the new linear layer and the final transformer layer were trained. The learning rate was again maintained at $3\times 10^{-4}$ within the AdamW optimiser, with a binary cross entropy loss function. The model was fine-tuned over 5000 iterations, with a batch size of 128.
To train and evaluate the fine-tuning of the model for beat detection, the “MIMIC PERform” training and test datasets were used [18]. Both datasets consist of real-world clinical PPG and ECG recordings at a sample frequency of 125Hz from 200 subjects, for 10 minutes per subject. The training dataset was used to train and validate, and the unseen test dataset was then used to evaluate performance of the trained model. The PPG signals were band-pass filtered between 1 and 30 Hz to remove low frequency variations, and ECG between 4 and 60Hz to isolate the higher frequency QRS complex and remove P and T waves. Labels for peaks were determined by using the inbuilt MATLAB findpeaks function on the filtered reference ECG data. Peaks, and the points immediately neighbouring the peaks, were labelled as 1. All other points were labeled as zero. This signal was then aligned with the filtered PPG signal to the point of maximal cross correlation, to ensure peak labels that mapped to the PPG peaks. Both the reference signal and the PPG signal were resampled to 50Hz, and the PPG signal was then tokenised as per previous examples. In the reference signal, the difference in the time between each peak should be minimal. When the standard deviation of this difference was greater than 0.3 seconds, a sign that the reference signal was of poor quality, the subject was excluded from the training data.
When testing, the criterion outlined by Chartlon et al. [18] was followed to allow for comparability of results. Poor quality labels were excluded from the test set by removing reference segments which had no beats for 2 seconds or more. Input segments with constant values for more than 0.2 seconds, an indication of clipping and thus data loss, were also excluded from analysis. Output peaks were determined by applying findpeaks to the output of the model, with a minimum height of 0.015, a maximum peak width of 0.4 seconds, and a minimum distance between peaks of 0.28 seconds (corresponding to a heart rate of 214 beats per minute). If a predicted beat was within 0.15 seconds of a reference beat, it was deemed to be correct. Results were then analysed in terms of the median F1 score (%) across test subjects.
It should be noted that the fine-tuned model produces continuous values between 0 and 1, for each position in the context, which are a measure of its confidence that a beat is present. In our analysis, we also investigated the utility of using these beat confidence values as a measure of signal quality.